OpenStudy (anonymous):

The tenth term of (x+1)^20 is 167,960x^11. Write the twelfth term.( hint: The eleventh term is the middle term)

OpenStudy (anonymous):

@amistre64 @SolomonZelman @yrelhan4

OpenStudy (anonymous):

\[\left( x+a \right)^n ,Tr+1=nCr ~x ^{n-r}~a ^{r}\] put n=20,r+1=12,r=12-1=11 a=1

OpenStudy (anonymous):

\[nCr=20C11=20C(20-11)=20C9\]

OpenStudy (anonymous):

I'm supposed to use Pascal's Triangle to figure this out... but i don't know how to apply it to this problem.

OpenStudy (anonymous):

\[\frac{ 20*19*18*17*16*15*14*13*12 }{ 9*8*7*6*5*4*3*2*1 }=?\]

OpenStudy (anonymous):

how did you get that? this is confusing :/

OpenStudy (anonymous):



OpenStudy (anonymous):

\[x ^{20-11}=x^9,a ^{11}=1^{11}=1\] so term is 197960 x^9

OpenStudy (anonymous):

in the previous solution i have used binomial theorem.  Term is \[x ^{11-2}=x^9\] power of x goes on decreasing.

Parth (parthkohli):

Hi, are you supposed to use the Pascal's Triangle?

OpenStudy (anonymous):

yes, that's what they want me to usee.

Parth (parthkohli):

If so, refer the 20th row. According to the hint, the eleventh term is the middle term, so you're looking for the number in the middle number in that row.

OpenStudy (anonymous):

how do I find the 20th row?

Parth (parthkohli):

It's really not viable to see Pascal's here. It's hard to write everything down to 20 rows. :P

Parth (parthkohli):

Do you know where the coefficients in Pascal's Triangle come from?

Parth (parthkohli):

(That's exactly what surjithayer showed you.)

OpenStudy (anonymous):

each of the numbers is the sum of the two numbers above it?

OpenStudy (anonymous):

Here are all the terms http://www.wolframalpha.com/input/?i=expand+%28x%2B1%29^20

Parth (parthkohli):

Yes, but I'm talking about another way. Do you know about permutations and combinations?

OpenStudy (anonymous):

no, I don't.

Parth (parthkohli):

Hmm, then you'll have to refer to a huge version of Pascal's. Let me search for it.

OpenStudy (anonymous):

yeah I'm learning permutations pretty much at the end of my algebra 2 textbook.

Parth (parthkohli):

Here, I found the twentieth row. {1, 20, 190, 1140, 4845, 15504, 38760, 77520, 125970, 167960, 184756, 167960, 125970, 77520, 38760, 15504, 4845, 1140, 190, 20, 1} Find the eleventh number in that.

Parth (parthkohli):

@eliassaab They haven't done combinatorics yet. :(

Parth (parthkohli):

Oh wait a second, they need to find the twelfth term.

Parth (parthkohli):

Then the twelfth number in that list of numbers is the coeff.

OpenStudy (anonymous):

okay now how would i know what power x would be?

OpenStudy (anonymous):

oh I know how, it starts out to the twentieth power and then gets smaller.

OpenStudy (ranga):

Polynomials are written from highest power of x to the lowest and so I would say x^9.

OpenStudy (anonymous):

so yeah x^9.

Parth (parthkohli):

Let's try to give you a simple example.\[(x+1)^2 = x^2 + 2x + 1\]The first term has power \(n\), the second has power \(n-1\) ... the \(k\)th term has power \(n - k+1\).

OpenStudy (anonymous):

thanks everyone for the help... thanks for the twentieth row of Pascal's @ParthKohli

Parth (parthkohli):

Hence, the twelfth term has power \(20 - 12 + 1 = 9\).

Parth (parthkohli):

No problem!

OpenStudy (anonymous):

and I just read on in my book and this was the next section of it!

Parth (parthkohli):

Ah, cool.